# The Population Resemblance Statistic: A Chi-Square Measure of Fit for Banking

### v3

@@ 5 Banking applications and performance analysis
If we considered an alternative setup for the methodology, say $c=0.9$, $M=1.5$ and $\alpha_{1}=0.2$, $\alpha_{2}=0.1$, the only differently assigned statuses (now $\mathbbm{g}$ i.s.o. $\mathbbm{a}$) would occur at $t_{1}$ of $(500,10)$, $t_{2}$ of $(n,B)=(2,000,10)$; as well as a now $\mathbbm{a}$ status instead of $\mathbbm{r}$ at $t_{6}$ of $(10,000,20)$. Here, the “partial discrepancy” ($\mathbbm{a}$) region is smaller, as evident from the choices of $M$ and larger values of $\alpha_{1}$ and $\alpha_{2}$. Nonetheless, there is a large degree of overlap between the two sets of results.
@@
A comparison with the discrete KS remains. As with the PSI(YN), the KS is based on a null hypothesis of exact equality, while the PRS includes a risk tolerance through $\delta$. The KS is also not distribution-free with respect to $\bm{p}_{0}$, a further drawback. As a first departure, observe Table 3: the PRS is more likely to indicate discrepancy ($\mathbbm{r}$, or $\mathbbm{a}$) than the KS. Further to this, it is well known that the power of the KS is small at small sample sizes (not unlike the PSI using YN critical values) possibly explaining the more frequent $\mathbbm{g}$ status where $n=50$, compared to the larger sample sizes. In all other cases (i.e. Tables 4 and 5), the KS signals full discrepancy ($\mathbbm{r}$) almost always. This is a direct result of the substantial power of the KS at larger sample sizes, and by its design, that any shift away from the null will swiftly be detected. Perhaps, detecting even the slightest shifts so often might not be practically ideal to a risk practitioner. Utilizing the PRS in these cases allows for a range of detection capabilities (see case $t_{1}$ and $t_{3}$ of $(n,B)=(10,000,20)$ with an $\mathbbm{a}$ and $\mathbbm{g}$ status, respectively).
@@
We conclude from this comparative real-world study that among the measures considered here, the PRS is universally competent at a range of sample sizes, including small sample sizes. A clear advantage of the PRS over the PSI and KS is the inclusion of the concept of $\delta$-resemblance and the tuning parameters $c$ and $M$ allowing the practitioner to calibrate the procedure to align with their risk tolerance. The PRS clearly indicates discrepancy sufficiently in smaller samples and often enough in larger samples. Unlike the KS, the PRS is sensitive to ranges of shifts over multiple risk categories and has easily obtainable critical values that are unique in small sample sizes.
@@ 6 Conclusion
Monitoring for changes in the population underlying a developed model is a common practice, especially in credit risk modeling. Over the years, several measures – most notably the PSI – have been proposed. However, limitations in these methods have spurred research into alternative approaches for assessing population resemblance. Our contribution, the PRS, utilizing the Pearson chi-square statistic and non-central chi-square distribution to address these gaps.
The main advantageous features are that the PRS accommodates sample-size dependent critical values and its explicit specification of risk tolerance. The PRS is statistically well-founded, performing reliably across a range of sample sizes, including in small samples. Unlike the discrete Kolmogorov-Smirnov test, the PRS is asymptotically distribution-free with respect to $\bm{p}_{0}$. Further, the PRS critical values are designed to account for acceptable levels of population shift, aligning with a composite null hypothesis framework. The risk tolerance $\delta$ explicitly incorporates the assumption limited population shift, distinguishing the PRS from the KS test and the PSI of Lewis and Yurdakul and Naranjo, (2020), which do not accommodate this structured flexibility. The tolerance parameter is incorporated through the concept of $\delta$-resemblance, a convenient way to communication population shift to practitioners in using the concept to derive business outcomes.

### v4

@@ 5 Banking applications and performance analysis
If we considered an alternative setup for the methodology, say $c=0.9$, $M=1.5$ and $\alpha_{1}=0.1$, $\alpha_{2}=0.2$, the only differently assigned statuses (now green i.s.o. amber) would occur at $t_{1}$ of $(500,10)$, $t_{2}$ of (n,B)=(2,000,10)(n,B)=(2{,}000,10); as well as an amber status instead of red at $t_{6}$ of (10,000,20)(10{,}000,20). Here, the amber region is smaller, as evident from the choices of $M$ and larger values of $\alpha_{1}$ and $\alpha_{2}$. Nonetheless, there is a large degree of overlap between the two sets of results.
@@
A comparison with the discrete KS remains. As with the PSI(YN), the KS is based on a null hypothesis of exact equality, while the PRS includes a risk tolerance through $\delta$. The KS is also not distribution-free with respect to $\bm{p}_{0}$, a further drawback. As a first departure, observe Table 3A: the PRS is more likely to indicate discrepancy (red or amber) than the KS. Further to this, it is well known that the power of the KS is small for small sample sizes (not unlike the PSI using YN critical values) possibly explaining the more frequent green status where $n=50$, compared to the larger sample sizes. In all other cases (Tables 3B, 3C and 4), the KS signals full discrepancy (red) almost always. This is a direct result of the substantial power of the KS at larger sample sizes, and by its design, that any shift away from the null will swiftly be detected. Perhaps, detecting even the slightest shifts so often might not be ideal to a risk practitioner. Utilizing the PRS in these cases allows for a range of detection capabilities (see case $t_{1}$ and $t_{3}$ of (n,B)=(10,000,20)(n,B)=(10{,}000,20) with an amber and green status, respectively).
@@
We conclude from this comparative real-world study that, among the measures considered here, the PRS is universally competent at a range of sample sizes, including small sample sizes. A clear advantage of the PRS over both the PSI and KS is the inclusion of the concept of $\delta$-resemblance and the tuning parameters $c$ and $M$ allowing the practitioner to calibrate the procedure to align with their risk appetite. The PRS clearly indicates discrepancy sufficiently in smaller samples and often enough in larger samples. Unlike the KS, the PRS is sensitive to ranges of shifts over multiple risk categories and has easily obtainable critical values that are unique in small sample sizes.
@@ 6 Conclusion
Monitoring for changes in the population underlying a developed model is a common practice, especially in credit risk modeling. Over the years, several measures — most notably the PSI — have been proposed. However, limitations in these methods have spurred research into alternative approaches for assessing population resemblance. Our contribution, the PRS, utilizes the Pearson chi-square statistic and non-central chi-square distribution to address these gaps.
The main advantageous features are that the PRS accommodates sample-size dependent critical values and its explicit specification of risk tolerance. The PRS is statistically well-founded, performing reliably across a range of sample sizes, including in small samples. Unlike the discrete Kolmogorov-Smirnov test, the PRS is asymptotically distribution-free with respect to $\bm{p}_{0}$. Further, the PRS critical values are designed to account for acceptable levels of population shift, aligning with a composite null hypothesis framework. The risk tolerance $\delta$ explicitly incorporates an assumption of limited population shift, distinguishing the PRS from the KS test and the PSI of Lewis and [yurdakul2020statistical], which do not accommodate this structured flexibility. The tolerance parameter is incorporated through the concept of $\delta$-resemblance, a convenient way to communication population shift to practitioners in using the concept to derive business outcomes.
@@
To conclude, we gratefully acknowledge the reviewers for their insightful and constructive feedback, which significantly strengthened the manuscript. Their comments helped clarify the composite null hypothesis framework, refine the exposition of the proposed procedure, suggest connections to stochastic ordering, and led to a more general and rigorous formulation and proof of Proposition 2.